Schedule of cash payments for a service company

Horizon Financial Inc. was organized on February 28. Projected selling and administrative expenses for each of the first three months of operations are as follows:

March $103,700
April 97,500
May 88,700

Depreciation, insurance, and property taxes represent $22,000 of the estimated monthly expenses. The annual insurance premium was paid on February 28, and property taxes for the year will be paid in June. 67% of the remainder of the expenses are expected to be paid in the month in which they are incurred, with the balance to be paid in the following month.

Prepare a schedule of cash payments for selling and administrative expenses for March, April, and May.

Excel Learning Systems Inc.
Schedule of Cash Payments for Selling and Administrative Expenses
For the Three Months Ending May 31
March April May
March expenses:
Paid in March $
Paid in April $
April expenses:
Paid in April $
Paid in May $
May expenses:
Paid in May $
Total cash payments $ $ $
Answer #1
Excel Learning Systems Inc.
Schedule of Cash Payments for Selling and Administrative Expenses
For the Three Months Ending May 31
March April May
March expenses:
Paid in March 54739
Paid in April 26961
April expenses:
Paid in April 50585
Paid in May 24915
May expenses:
Paid in May 44689
Total cash payments 54739 77546 69604
Formulas:
March expenses:
Paid in March =(103700-22000)*67%
Paid in April =(103700-22000)*33%
April expenses:
Paid in April =(97500-22000)*67%
Paid in May =(97500-22000)*33%
May expenses:
Paid in May =(88700-22000)*67%